# Language Modeling Is Compression

### v1

@@ 1 Introduction
Arithmetic coding, in particular, is known to be optimal in terms of coding length, meaning that the overall compression performance depends on the capabilities of the probabilistic model (Fig. 1). Incidentally, in recent years, large pre-trained Transformers (Vaswani et al., 2017), so-called foundation models (Bommasani et al., 2021), have proven to be highly successful across a wide range of predictive tasks (Bubeck et al., 2023; Rae et al., 2021) and are thus promising candidates for use with arithmetic coding. Indeed, Transformer-based compression with arithmetic coding has produced state-of-the-art results both in the online (Bellard, 2021; Mao et al., 2022) and offline settings (Valmeekam et al., 2023). In the online setting, a pseudo-randomly initialized model is directly trained on the stream of data that is to be compressed, while the offline setting, which we consider in our work, trains the model on an external dataset before employing it to compress a (potentially different) data stream. Consequently, offline compression is performed in-context, with a fixed set of model parameters. Transformers have demonstrated impressive in-context learning abilities (Laskin et al., 2023; Brown et al., 2020; Wei et al., 2022; Genewein et al., 2023), which renders them ideally suited for offline compression. However, as we will discuss in this work, Transformers are actually trained to compress well, and therefore must have good in-context learning abilities.
@@
We advocate for using (lossless) compression to study foundation models. To that end, we conduct an extensive empirical investigation of the offline (in-context) compression capabilities of large language models, with the rationale that they have recently become readily available (Hoffmann et al., 2022; Touvron et al., 2023) and can thus be used for compression without the training overhead. We empirically demonstrate that these models, while (meta-)trained primarily on text, also achieve state-of-the-art compression rates across different data modalities, using their context to condition a general-purpose compressor to excel at a particular task. Moreover, we shed new light on scaling laws (Kaplan et al., 2020), showing that they also hold true for compression but that measuring the compression rates instead of the log loss adds a twist: Scaling beyond a certain point will deteriorate the compression performance since the model parameters need to be accounted for in the compressed output. Finally, we advocate for framing (self-supervised) prediction through the lens of compression as it encompasses generalization: a model that compresses well generalizes well (Hutter, 2006).
Figure 1: Arithmetic encoding of the sequence ‘AIXI’ with a probabilistic (language) model $P$ (both in blue) resulting in the binary code ‘0101001’ (in green). Arithmetic coding compresses data by assigning unique intervals to symbols based on the probabilities assigned by $P$. It progressively refines these intervals to output compressed bits, which represent the original message. To decode, arithmetic coding initializes an interval based on the received compressed bits. It iteratively matches intervals with symbols using the probabilities given by $P$ to reconstruct the original message.
@@
We make the following contributions:
  • We empirically investigate the lossless compression capabilities of foundation models. To that end, we review how to compress with predictive models via arithmetic coding and call attention to the connection between current language modeling research and compression.
  • We show that foundation models, trained primarily on text, are general-purpose compressors due to their in-context learning abilities. For example, Chinchilla 70B achieves compression rates of 43.4% on ImageNet patches and 16.4% on LibriSpeech samples, beating domain-specific compressors like PNG (58.5%) or FLAC (30.3%), respectively.
  • We provide a novel view on scaling laws, showing that the dataset size provides a hard limit on model size in terms of compression performance and that scaling is not a silver bullet.
@@ 2 Background
Above we discussed optimal (arithmetic) coding with respect to data sampled from a fixed distribution $\rho$. In contrast, universal (optimal) source coding with respect to all computable sampling distributions can, in theory, be achieved by choosing $\ell_{c}(x_{1:n})$ as the Kolmogorov complexity of $x_{1:n}$ (Kolmogorov, 1998; Li & Vitányi, 2019). For this choice, the conditional distribution described above is universally optimal over $x_{<i}$, recovering the Solomonoff predictor (Solomonoff, 1964a, b; Rathmanner & Hutter, 2011). The Solomonoff predictor is a Bayesian mixture of all predictors that can be programmed in a chosen Turing-complete programming language. More precisely, for a predictor $q$ of program-length $\ell_{c}(q)$ bits, the Solomonoff predictor assigns a prior weight of $2^{-\ell_{c}(q)}$ to predictor $q$. That is, if $\mathcal{Q}$ is the set of all predictors that can be programmed and computed, the Solomonoff predictor assigns probability $S(x_{1:n})=\sum_{q\in{\mathcal{Q}}}2^{-\ell_{c}(q)}q(x_{1:n})$ to a sequence $x_{1:n}$, if every predictor $q$ assigns that sequence probability $q(x_{1:n})$. Therefore, $S(x_{1:n})\geq 2^{-\ell_{c}(q)}q(x_{1:n})$ for all $q\in\mathcal{Q}$, and thus $-\log_{2}S(x_{1:n})\leq-\log_{2}q(x_{1:n})+\ell_{c}(q)$. Observe that $\ell_{c}(q)$ is a constant of $q$ that is independent of the sequence length. Therefore, compressing optimally is equivalent to predicting optimally and vice versa (Hutter, 2005).
@@ 3 Experimental Evaluation
We now present our evaluation of the (in-context) compression capabilities of foundation models.
@@
We compare our arithmetic coding-based language model compressors to two competitive general-purpose lossless compressors: gzip (Deutsch, 1996) and its improvement LZMA2 (Pavlov, 2019), used by the 7zip software. Both are based on Huffman coding (Huffman, 1952) and the Lempel-Ziv-Welch algorithm (Welch, 1984). We also consider specialized lossless compressors for image and audio data, i.e., PNG (Boutell, 1997) and FLAC (Coalson, 2008), respectively. Finally, we evaluate two types of language models (of different sizes) with arithmetic coding: vanilla decoder-only Transformers (Vaswani et al., 2017), which we pretrain on the enwik8 dataset, and pretrained Chinchilla-like foundation models (Hoffmann et al., 2022).
@@ 3.1 Datasets
[…]
@@
LibriSpeech (Panayotov et al., 2015) is a corpus of approximately $1000$ hours of 16kHz English speech. The data is derived from audiobooks from the LibriVox project and has been carefully segmented and aligned. We chunk the samples into batches of 2048 bytes and gather 488 821488821488\,821488 821 such chunks into dataset of size 1 GB.
3.2 Comparing Compression Rates
Table 1: Compression rates (compressed size / raw size) on different datatsets (lower is better). The raw compression rate does not take the parameter size into account for the Transformer and Chinchilla models, while the adjusted compression rate considers the parameter size part of the compressed size. All datasets are of raw size 1GB. Random data is used as a baseline and should not be compressible. Transformer and Chinchilla are predictive models, which we use with arithmetic coding to obtain lossless compressors. We train the Transformer models from scratch on enwik8, while the Chinchilla models are pretrained on large text datasets. Transformers trained on enwik overfit to that data modality, while Chinchilla models are good compressors for various data types.
@@ 3.2 Comparing Compression Rates
Chunk Size Compressor enwik9 ImageNet LibriSpeech Random enwik9 ImageNet LibriSpeech Random
@@
Table 1 shows the compression rates for all compressors and datasets. We show both the raw compression rate, which does not take the model size (in bytes) into account, as well as the adjusted rate, which does. The size of the Python program for classical compressors is very small (a few kilobytes at most) and thus barely affects the compression rate. In contrast, language models suffer a huge loss in compression rate due to their large size, which cannot be offset when compressing only 1GB of data. We encode each neural network parameter with 2 bytes, using a float16 representation since quantizing weights to this level does not significantly affect performance (Tao et al., 2022) and is standard for model inference. Note that further compressing the float16 parameters using classical compressors does not significantly reduce their size (we obtained rates of 92.2% and 89.1% on a 38M parameter Transformer with gzip and LZMA2, respectively). Also, recall that we only consider the offline setting, which computes the adjusted compression rate using a two-part code (i.e., it adds the model size to the $\log$-loss of the data). In contrast, prequential (online) coding would provide an alternative view on adjusted compression by computing the adjusted compression rate as the $\log$-loss plus the size of the training script (not the model parameters). According to prior work, prequential coding leads to better compression with overparametrized neural networks (Blier & Ollivier, 2018), however, it requires training the model online (which reduces performance and cannot be performed with foundation models) both during encoding and decoding (which is very costly for our models).
Figure 2: Adjusted compression rates (compressed size / raw size) for Transformers of different sizes, trained on enwik8 and evaluated on enwik (both axes are logarithmic). Here, the compressed size does not only consider the size of the compressed output (roughly equal to the $\log$-loss) but also the model size, which causes all curves to increase at some point. Every dataset gives rise to an optimal model size, with a good trade-off between performance (the size of the compressed data) and cost of the model (the number of parameters). The larger the dataset, the more parameters we can afford.
@@
A lossless compressor induces an injective function over bit sequences, meaning that we cannot compress all sequences equally well (by the pigeonhole principle). Consequently, in practice, compressors are often tailored to a particular setting, e.g., FLAC for audio or PNG for images, and thus fail to compress other data modalities well (see Table 1). In contrast, general-purpose compressors, such as gzip, offer good performance on a wide range of data sources. Surprisingly, Chinchilla models, while trained primarily on text, also appear to be general-purpose compressors, as they outperform all other compressors, even on image and audio data (see Table 1). Note that Chinchilla models have not been trained on this kind of data according to Appendix A. of Hoffmann et al. (2022), which states that the training dataset consists of a mix of internet text data (Wikipedia, websites, github) and books. However, it is still possible (but unlikely) that some images or audio samples were encoded into text on some websites. Thus, Chinchilla models achieve their impressive compression performance by conditioning a (meta-)trained model to a particular task at hand via in-context learning (Genewein et al., 2023). In contrast, smaller Transformers, trained manually on enwik8, only achieve good compression rates on similar Wikipedia data, i.e., enwik9. However, larger models’ stronger in-context compression (or in-context learning) comes at a price: the number of parameters, which has to be offset with increasingly large data sources when computing the adjusted compression rate (see Section 3.3). Finally, note that since Chinchilla has been trained on Wikipedia, the enwik9 results are in-distribution.
@@ 3.3 Optimal Model-Dataset Size Tradeoff
Fig. 2 visualizes the adjusted compression rate for vanilla Transformers of different model sizes for the enwik datasets. We observe that larger models achieve better compression rates on larger datasets, thus justifying recent trends in model scaling (Kaplan et al., 2020). However, they achieve worse rates on smaller datasets, indicating that scaling laws are, in fact, dependent on the size of the test set. That is, for each dataset, the model sizes reach a critical point, after which the adjusted compression rate starts to increase again since the number of parameters is too big compared to the size of the dataset. Note that we evaluate offline compression, i.e., we do not necessarily compress the data the model was trained on, meaning that the results on enwik7 and enwik8 are in-distribution, while the enwik9 results are out-of-distribution. Nevertheless, larger models still achieve better compression rates on enwik9 than enwik8, illustrating the benefits of scaling.
@@ 3.4 Compressors as Generative Models
In Section 2, we discussed how any compressor can be employed as a sequence prediction model. Concretely, for compressor $c$, we sample the next byte according to the distribution $\hat{\rho}(x_{i}\mid x_{<i})\sim 2^{\ell_{c}(x_{<i})-\ell_{c}(x_{<i}x_{i})}$, i.e., we compute the length $\ell_{c}$ of the compressed sequence $c(x_{<i}b)$ for all possible $b\in\mathcal{X}$. Thus, if a byte $b$ leads to a particularly short compressed sequence (when concatenated with $x_{<i}$), it will have a higher probability of being sampled next. Note that any constant in the length function (e.g., the header for classical compressors) disappears when we normalize the distribution.
Since generic compressors have a low intrinsic bias, sampling data without conditioning does not yield interesting results as it looks random. Thus, we condition the compressors on part of an existing sequence (1948 bytes for enwik9, half of the sample for ImageNet and LibriSpeech) and generate the remaining bytes with the compression-based generative model. We compare the generative performance of gzip and Chinchilla 70B across all three data modalities in Figs. 3, 5 and 4 for text, image, and audio data, respectively. In general, generative models can be evaluated using one of two ways: sampling the next byte $\hat{\rho}(x_{i}\mid x_{<i})$ (i) using teacher forcing, i.e., conditioning on the true subsequence $x_{<i}$, or (ii) via autoregressive sampling, i.e., conditioning on the model’s previous outputs. The latter induces a distribution shift, and with it undesired side effects (Ortega et al., 2021), but is standard and thus what we choose to visualize.
Context Text (1948 Bytes)
Ground Truth (100 Bytes)
gzip Samples (100 Bytes)
Chinchilla 70B Samples (100 bytes)
Figure 3: Compression-based generation for text data. We condition gzip and Chinchilla on a context text of size 1948 bytes (from enwik9) and then sample 100 bytes ($N$ tokens) autoregressively. Since Chinchilla employs a tokenizer, the sampled sequences will contain $N$ tokens, which do not necessarily decode to 100 bytes. Chinchilla’s predictions are significantly more coherent than gzip’s.
(a) Original spectrogram
(b) gzip
(c) Chinchilla
Figure 4: Compression-based generation for audio data. We condition gzip and Chinchilla on the first 1024 bytes of the base sequence (from LibriSpeech) and then sample the remaining 1024 bytes autoregressively. Chinchilla predictions exhibit a typical “loop” pattern of autoregressive generation.

### v2

@@ 1 Introduction
Arithmetic coding, in particular, is known to be optimal in terms of coding length, meaning that the overall compression performance depends on the capabilities of the probabilistic model (see Fig. 1 for an overview of arithmetic coding). Incidentally, in recent years, large pre-trained Transformers (Vaswani et al., 2017), so-called foundation models (Bommasani et al., 2021), have proven to be highly successful across a wide range of predictive tasks (Bubeck et al., 2023; Rae et al., 2021) and are thus promising candidates for use with arithmetic coding. Indeed, Transformer-based compression with arithmetic coding has produced state-of-the-art results both in the online (Bellard, 2021; Mao et al., 2022) and offline settings (Valmeekam et al., 2023). In the online setting, a pseudo-randomly initialized model is directly trained on the stream of data that is to be compressed, while the offline setting, which we consider in our work, trains the model on an external dataset before employing it to compress a (potentially different) data stream. Consequently, offline compression is performed in-context, with a fixed set of model parameters. Transformers have demonstrated impressive in-context learning abilities (Laskin et al., 2023; Brown et al., 2020; Wei et al., 2022; Genewein et al., 2023) and are thus ideally suited for offline compression.
@@
We advocate for using (lossless) compression to study foundation models. To that end, we conduct an extensive empirical investigation of the offline (in-context) compression capabilities of large language models, with the rationale that they have become readily available (Touvron et al., 2023a; b) and can thus be used for compression without the training overhead. We empirically demonstrate that these models, while (meta-)trained primarily on text, achieve competitive compression rates across different data modalities, outperforming domain-specific standard compressors (not accounting for model parameter size). Moreover, we shed new light on scaling laws (Kaplan et al., 2020), showing that they also hold true for compression but that measuring the adjusted compression rates instead of the log loss adds a twist: Scaling beyond a certain point will deteriorate the compression performance since the model parameters need to be accounted for in the compressed output. Finally, we advocate for framing (self-supervised) prediction through the lens of compression as it encompasses generalization: a model that compresses well generalizes well (Hutter, 2006).
Figure 1: Arithmetic encoding of ‘AIXI’ with a probabilistic model $P$ (blue) resulting in the binary code ‘b0101010’ (green). We iteratively divide the real interval $I=[0,1)$ according to the model’s (conditional) probabilities and select the sub-interval corresponding to the observed symbol (e.g., $I=[0,0.45)$ for $P(A)$). We further refine $I$ for each input symbol (indicated by the arrows), e.g., $I=[0.09,0.36)$ for $P(I|A)$. To determine the encoded output, we iteratively split $[0,1)$ in half and assign a binary code to each sub-interval (shaded red areas). At every step we can output the binary code if $I$ is fully contained in the corresponding binary interval (e.g., ‘b0’ for ‘A’, but not for ‘AI’ as it could be ‘b00’ or ‘b01’). At the end of the input, the code is ‘b0101’, which cannot be uniquely decoded ($P(A|AIX)$, $P(I|AIX)$, $P(X|AIX)$ all overlap with ‘b0101’). Thus, we further refine the binary code until its binary interval is fully contained in $I$ (all calculations in Appendix A).
@@
We empirically study the lossless compression capabilities of foundation models:
  • We review how to compress with predictive models via arithmetic coding and call attention to the connection between current language modeling research and compression.
  • We show that large language models achieve impressive compression rates (disregarding model parameter size) on modalities other than text. For example, Chinchilla 70B achieves compression rates of 43.4% on ImageNet patches and 16.4% on LibriSpeech samples, beating domain-specific compressors like PNG (58.5%) or FLAC (30.3%), respectively.
  • We revisit scaling laws, showing that the dataset size provides a hard limit on model size in terms of compression performance and that model scaling is not a silver bullet.
@@ 2 Background
Above we discussed optimal (arithmetic) coding with respect to data sampled from a fixed distribution $\rho$. In contrast, universal (optimal) source coding with respect to all computable sampling distributions can, in theory, be achieved by choosing $\ell_{c}(x_{1:n})$ as the Kolmogorov complexity of $x_{1:n}$ (Kolmogorov, 1998; Li & Vitányi, 2019). For this choice, the conditional distribution described above is universally optimal over $x_{<i}$, recovering the Solomonoff predictor (Solomonoff, 1964a; b; Rathmanner & Hutter, 2011). The Solomonoff predictor is a Bayesian mixture of all predictors that can be programmed in a chosen Turing-complete programming language. More precisely, for a predictor $q$ of program-length $\ell_{c}(q)$ bits, the Solomonoff predictor assigns a prior weight of $2^{-\ell_{c}(q)}$ to predictor $q$. That is, if $\mathcal{Q}$ is the set of all predictors that can be programmed and computed, the Solomonoff predictor assigns probability $S(x_{1:n})=\sum_{q\in{\mathcal{Q}}}2^{-\ell_{c}(q)}q(x_{1:n})$ to a sequence $x_{1:n}$. Therefore, $S(x_{1:n})\geq 2^{-\ell_{c}(q)}q(x_{1:n})$ for all $q\in\mathcal{Q}$, and thus $-\log_{2}S(x_{1:n})\leq-\log_{2}q(x_{1:n})+\ell_{c}(q)$. Observe that $\ell_{c}(q)$ is a constant of $q$ that is independent of the sequence length. Therefore, compressing optimally is equivalent to predicting optimally and vice versa (Hutter, 2005).
@@ 3 Experimental Evaluation
Here, we evaluate foundation models’ (in-context) compression capabilities (details in Appendix B and code at https://github.com/google-deepmind/language_modeling_is_compression).
@@
We compare our arithmetic coding-based language model compressors to two competitive general-purpose lossless compressors: gzip (Deutsch, 1996) and its improvement LZMA2 (Pavlov, 2019), used by the 7zip software. Both are based on Huffman coding (Huffman, 1952) and the Lempel-Ziv-Welch algorithm (Welch, 1984). We also consider specialized lossless compressors for image and audio data, i.e., PNG (Boutell, 1997) and FLAC (Coalson, 2008), respectively. Finally, we evaluate two types of language models (of different sizes) with arithmetic coding: vanilla decoder-only Transformers (Vaswani et al., 2017), which we train on the enwik8 dataset, and the pretrained Llama 2 (Touvron et al., 2023b) and Chinchilla (Hoffmann et al., 2022).
@@ 3.1 Datasets
A key question is how to reconcile the different context lengths $C$ of the compressors we consider. Transformers are restricted to short contexts (2048 “tokens”, coded over 1 byte for our trained transformers, and 4 bytes for the pretrained models), while gzip uses a maximum context of 32 kilobytes, and LZMA2 has a virtually “infinite” context length. Having a longer context allows a compressor to exploit more sequential dependencies to achieve a better compression rate. For compressors with finite contexts, there are two approaches to compress sequences that are longer than the context length: (i) slide the compressor byte by byte, thus always processing a history of the previous $C-1$ bytes when compressing a new byte, and (ii) chunk the data stream into $S$ sequences of $C$ bytes and evaluate the in-context compression (without any history) averaged across batches. For Transformers, we consider the latter approach since sliding would increase their (already very long) running time by a factor of $S$. Therefore, we chunk all datasets into sequences of $2048$ bytes and feed them to the compressors one-by-one. However, since classical compressors usually include a header in their compressed output, which can be larger than the compressed data in some cases, we only count it once for all batches. Moreover, since chunking deteriorates the performance of classical compressors, which have context lengths $C\gg 2048$, we also report their compression rates on the unchunked datasets. We consider the following datasets:
@@
LibriSpeech (Panayotov et al., 2015) contains roughly $1000$ hours of 16kHz English speech data derived from audiobooks of the LibriVox project that has been segmented and aligned. We chunk the samples into 2048 bytes and gather 488 821488821488\,821488 821 such chunks into a dataset of size 1 GB.
Table 1: Compression rates (compressed size / raw size) on different datatsets (lower is better). The raw compression rate does not take the parameter size into account for the neural models, while the adjusted compression rate considers the parameter size part of the compressed size. All datasets are of size 1GB. Random data is used as a baseline and should not be compressible. Transformer, Llama 2, and Chinchilla are predictive models, which we use with arithmetic coding to obtain lossless compressors. We train the Transformer models from scratch on enwik8, while the Chinchilla models are pretrained on large text datasets. Transformers trained on enwik overfit to that data modality, while large language models are good compressors for various data types.
@@
Chunk Compressor enwik9 ImageNet LibriSpeech Random enwik9 ImageNet LibriSpeech Random
@@
3.2 Comparing Compression Rates
Table 1 shows the compression rates for all compressors and datasets. We show both the raw compression rate, which does not take the model size (in bytes) into account, as well as the adjusted rate, which does. The size of the Python program for classical compressors is very small (a few kilobytes at most) and thus barely affects the compression rate. In contrast, language models suffer a huge loss in compression rate due to their large size, which cannot be offset when compressing only 1GB of data. We encode each neural network parameter with 2 bytes, using a float16 representation since quantizing weights to this level does not significantly affect performance (Tao et al., 2022) and is standard for model inference. Note that further compressing the float16 parameters using classical compressors does not significantly reduce their size (we obtained rates of 92.2% and 89.1% on a 38M parameter Transformer with gzip and LZMA2, respectively). We only consider the offline setting, which computes the adjusted compression rate using a two-part code (i.e., it adds the model size to the $\log$-loss of the data). In contrast, prequential (online) coding would provide an alternative view on adjusted compression by computing the adjusted compression rate as the $\log$-loss plus the size of the training script (not the model parameters). Prequential coding leads to better compression with overparametrized neural networks (Blier & Ollivier, 2018), but it requires training the model online both during encoding and decoding (which is very costly for our models).
@@ 3.2 Comparing Compression Rates
A lossless compressor induces an injective function over bit sequences, meaning that we cannot compress all sequences equally well (by the pigeonhole principle). Consequently, in practice, compressors are often tailored to a particular setting, e.g., FLAC for audio or PNG for images, and thus fail to compress other data modalities well (see Table 1). In contrast, general-purpose compressors, such as gzip, offer good performance on a wide range of data sources. Surprisingly, large language models, while trained primarily on text, also appear to be general-purpose compressors, as they outperform all other compressors, even on image and audio data (see Table 1). Note that these models have not been trained on this kind of data: for Chinchilla, Appendix A. of Hoffmann et al. (2022) states that the training dataset consists of a mix of internet text data (Wikipedia, websites, github) and books. However, it is still possible (but unlikely) that some images or audio samples were encoded into text on some websites. Thus, pretrained models achieve their impressive compression performance by conditioning a (meta-)trained model to a particular task at hand via in-context learning (Genewein et al., 2023). In contrast, smaller Transformers, trained manually on enwik8, only achieve good compression rates on similar Wikipedia data, i.e., enwik9. However, larger models’ stronger in-context compression (or in-context learning) comes at a price: the number of parameters, which has to be offset with increasingly large data sources when computing the adjusted compression rate (see Section 3.3).
[Uncaptioned image] Figure 2: Adjusted compression rates (compressed size / raw size) for Transformers of different sizes, trained on enwik8 and evaluated on enwik (both axes are logarithmic). Here, the compressed size does not only consider the size of the compressed output (roughly equal to the $\log$-loss) but also the model size, which causes all curves to increase at some point. Every dataset gives rise to an optimal model size, with a good trade-off between performance (the size of the compressed data) and cost of the model (the number of parameters). The larger the dataset, the more parameters we can afford.
Table 2: Raw compression rates (compressed size / raw size) on enwik9 for Transformers trained on enwik8 with different tokenizers, ASCII and byte-pair encoding (BPE), with various vocabulary sizes. Transformers compress better with simpler tokenizers. However, larger vocabulary sizes reduce the length of the sequence more, meaning more information can be packed into the context. Raw Compression Rate (%) Tokenizer 200K 6.4M 38M ASCII 22.9 13.6 6.4 BPE 1K 25.4 14.8 6.9 BPE 2K 25.6 15.7 7.4 BPE 5K 23.1 17.1 8.7 BPE 10K 21.3 17.0 8.9 BPE 20K 19.3 16.4 9.0
@@ 3.3 Optimal Model-Dataset Size Tradeoff
Fig. 2 visualizes the adjusted compression rate for vanilla Transformers of different sizes for enwik. We observe that larger models achieve better compression rates on larger datasets, justifying recent trends in model scaling (Kaplan et al., 2020). However, they achieve worse rates on smaller datasets, indicating that scaling laws are, in fact, dependent on the size of the test set. That is, for each dataset, the model sizes reach a critical point, after which the adjusted compression rate starts to increase again as the number of parameters overweighs the size of the dataset. Note that we evaluate offline compression, i.e., we do not necessarily compress the data the model was trained on, meaning that the results on enwik7 and enwik8 are in-distribution, while enwik9 is (partially) out-of-distribution.
@@ 3.4 Compressors as Generative Models
In Section 3.2 we showed that any predictor can be employed as a compressor. Here, following Section 2, we empirically demonstrate the opposite direction, i.e., that compressors can be used as a sequence prediction model, establishing our main claim that “language modeling is compression”. We compute the length $\ell_{c}$ of the compressed sequence $c(x_{<i}b)$ for all possible $b\in\mathcal{X}$ to get the probabilities $\hat{\rho}(b\mid x_{<i})=2^{\ell_{c}(x_{<i})-\ell_{c}(x_{<i}b)}$. This can straightforwardly be extended to sampling a whole continuation autoregressively by appending the last output to the sequence and iterating.
Theoretically, there is no strong guarantee that a good compression rate leads to “good” autoregressive samples. However, empirically it has been shown that better sequence prediction (i.e., lower $\log$-loss) often leads to better generation (Rae et al., 2021; Brown et al., 2020). Nevertheless, in autoregressive sampling small errors often accumulate, which can lead to samples that diverge from the ground-truth distribution. Also, this standard sampling technique only looks one step into the future, and can be biased: gzip, for instance, builds an internal dictionary of ’tokens’, which will be compressed using their indexes. Extending the sequence $x_{<i}$ with one of these tokens will lead to a good compression rate, but will be omitted as it can be longer than one byte. Our neural models do not suffer such bias as they are trained to predict one step ahead with the cross-entropy loss.
We compare the generative capabilities of gzip and Chinchilla 70B on images in Fig. 3. Each image is a sampled from ImageNet with height 290 and width 500. For each row in the image, we condition the model on the first 250 pixels and autoregressively generate the remaining 250 pixels, treating different rows as independent of each other (an oversimplification w.r.t. natural image statistics). We use the same byte conversions and tokenization details as explained in Appendix B. Chinchilla 70B shows signatures of visually appropriate continuations (judged qualitatively), which tend to degrade with increased sample length as more and more error accumulates. gzip produces much noisier completions. We compare the generative performance of gzip and Chinchilla (1B, 7B, and 70B) across all three data modalities in Figs. C.1, C.2 and C.3 for text, image, and audio data, respectively.